If 65.7 g of aspirin (C₉H₈O₄) are produced from 79.8 g of C₇H₆O₃, what is the percent yield from the reaction below?

C₇H₆O₃ (s) + C₄H₆O₃ (l) → C₉H₈O₄ (s) + HC₂H₃O₂ (aq).


1
Expert's answer
2021-11-15T10:56:30-0500

Taking "C=12,H=1, and \>O=16"


Molar weight of "C_7H_6O_3=(12\u00d77)+(1\u00d76)+(16\u00d73)"


"\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>=138g"


Molar weight of "C_9H_8O_4=(12\u00d79)+(1\u00d78)+(16\u00d74)"

"=180g"



Theoretically "138g" of "C_7H_6O_3" yields "180g" of Aspirin


"\therefore79.8g" of "C_7H_6O_3" should yield


"\frac{79.8}{138}×180=104.09g" of Aspirin



Actual yield "=65.7g"


"\therefore\%" Yield "=\frac{Actual\> Yield}{Theoretical \>Yield}×100\%"



"=\frac{65.7}{104.09}×100\%=63.118\%"